# Risk-Aware Real-Time Task Allocation for

### v1

@@ III-D Control Synthesis and Local Risk Value
In this subsection, we compute the local risk value and synthesize a controller for a given agent-specification pair. We assume a new specification $\phi$ is provided at time $k$ together with maximal risk $r_{\phi,\max}$. For any specifications $\psi$, we denote with $k_{\psi}$ the time at which the specification was assigned. We denote with $\mathcal{P}^{k}$ the set of all previously accepted specifications $\psi$ with $k_{\psi}<k$, together with the newly assigned specification $\phi$, i.e., $\phi\in\mathcal{P}^{k}$. The objective is to synthesize a controller $\bm{f}$ such that $\varphi=\wedge_{\psi\in\mathcal{P}^{k}}\lozenge_{[k_{\psi},k_{\psi}]}\psi$ is satisfied in accordance to the maximal risk of each conjunction element, i.e., $\mathbb{P}_{\bm{f}}(\psi,k_{\psi})\geq 1-r_{\psi,\max},\ \forall\psi\in\mathcal{P}^{k}$.
We are interested in implementing a tube-based MPC algorithm that at each time instance $k\in[0,N]$ recomputes the optimal nominal trajectory $\bm{z}_{[k,N]}$, the optimal nominal input $\bm{v}_{[k,N-1]}$, and the optimal risk levels $\bm{r}_{[k,N]}$. For this, we assume given $x(k)$ together with $\bm{z}_{[0,k]}$, $\bm{v}_{[0,k-1]}$, $\bm{\rho}_{[0,k]}$ and $\bm{r}_{[0,k]}$ computed at time $k-1$. Furthermore, for all $\psi\in\mathcal{P}^{k}$, we assume $\mathcal{H}_{\psi}^{k_{\psi}}\subseteq[0,N]$ without lose of generality.
@@
where $\Xi:=\{\bm{z}_{[k,N]},\bm{v}_{[k,N-1]},\bm{r}_{[k+1,N]},\bm{\rho}_{[k+1,N]},s(k)\}$, $J$ is either a linear or quadratic cost function, the constraints (9d)-(9e) are obtained from $\varphi=\wedge_{\psi\in\mathcal{P}^{k}}\lozenge_{[k_{\psi},k_{\psi}]}\psi$ and the constraints (9f)-(9g) are obtained from Prop. 3. From the above tube-based MPC, an (updated) control strategy is obtained from $\bm{v}_{[k,N-1]}$, and a local risk value $r_{\phi,mpc}$ is obtained from the right-hand side of (9g), i.e.,
@@
A solution to (9), by Prop. 3, ensures that for all $\psi\in\mathcal{P}^{k}$, $\mathbb{P}_{\bm{f}}(\psi,k_{\psi})\geq 1-r_{\psi,\max}$. Should there exist no solution to problem (9), the corresponding agent-specification pair is deemed infeasible. Consequently, in the single-agent case, we will then update the control strategy based only on measurement $x(k)$, as per constraint (9b). This can be achieved by considering the MPC problem (9) with $\mathcal{P}^{k}=\mathcal{P}^{k}\setminus\{\phi\}$. However, should also the measurement fail to produce a controller, the previously computed control strategy will be implemented instead.
IV Risk-Aware Multi-Agent Task Allocation
In this section, we consider the multi-agent case. More specifically, this section will primarily cover real-time task allocation for multi-agent systems. We refer to section III regarding control synthesis and local risk value computation for individual agent-specification pairs.
The section will be structured as follows. First, we will consider specification decomposition and heuristic agent-specification filtering. Second, we utilize auctioning, based on local risk values, to obtain a definitive agent-specification assignment. The section concludes with an implementation algorithm and theoretical analysis of the overall method.
IV-A Specification Decomposition
Let $\phi$ be a newly assigned STL specification with maximal risk $r_{\phi,\max}$, involving a distinct number of agents, $\nu\leq|\mathcal{I}|$. The objective is to decompose the specification $\phi$ into agent-level sub-specifications, $\phi^{j}$, $j\in\mathcal{I}$, and allocate individual maximal risks, $r_{\phi^{j},\max}$, in accordance to $r_{\phi,\max}$. First, we define $\mathrm{dec}(\phi,\nu)=\wedge_{j=1}^{\nu}\phi^{j}$ as the decomposition of $\phi$ into a conjunctive form, with each conjunct representing a sub-specification, $\phi^{j}$, the satisfaction of which depends on a single agent $j$, and $\nu$ being the number of agents needed to satisfy $\phi$. Second, we assign a maximal risk to $\phi^{j}$, following a union-bound argument, which requires $\sum_{j=1}^{\nu}r_{\phi^{j},\max}\geq r_{\phi,\max}$. In this paper, we select a uniform allocation of risks across sub-specifications, obtaining
$r_{\phi^{j},\max}=\frac{r_{\phi,\max}}{\nu},$ (11)
for all agents $j\in\{1,\ldots,\nu\}$.
The decomposition $\mathrm{dec}(\phi,\nu)$ acts as a transformation on $\phi$ returning a conjunctive formula, the satisfaction of which implies satisfaction of $\phi$. However, the converse is not necessarily true. There are a few methods [12, 13, 24] for splitting STL specifications into sub-specifications. In this paper, given a specification $\phi$, we convert Boolean and temporal operators to conjunctive forms based on the heuristics in [24] and decompose predicates following a similar approach to [12]. Therein, predicates are defined on general convex sets, whereas negated forms are not studied. Here, we focus on polytopic settings and allow negations. For our setting, we assume the following.
Assumption 2
In view of existing decomposition approaches, where sub-specifications are assigned to subteams of agents, Assumption 2 is restrictive. However, in the setting of polytopic predicates, we can produce agent-level sub-specifications by projecting the associated predicates onto the agents’ state space. Before showing this, we define the following.
Definition 4
Lemma 5
Let $\mu:=(h(x)\geq\bm{0})$ be a predicate, $H:=\{x\mid h(x)\geq\bm{0}\}$ be a polytope, with $x=[x_{1}^{\intercal}\;\ldots\;x_{\nu}^{\intercal}]^{\intercal}$, $x_{i}\in\mathbb{R}^{n}$, $i\in\{1,\ldots,\nu\}$, and $B_{\mathrm{in}}(H)\subseteq H$ and $B_{\mathrm{out}}(H)\supseteq H$ be the associated orthotopes inscribed in and circumscribing $H$, respectively. Then, it holds that (i) $\mathrm{dec}(\mu,\nu)=\wedge_{i=1}^{\nu}\mu_{\mathrm{in}}^{i}$ satisfies $\mu$, with $\mu_{\mathrm{in}}^{i}:=(x_{i}\in\mathrm{Proj}_{x_{i}}(B_{\mathrm{in}}(H))$, and (ii) $\mathrm{dec}(\neg\mu,\nu)=\wedge_{i=1}^{\nu}\neg\mu_{\mathrm{out}}^{i}$ satisfies $\neg\mu$, with $\mu_{\mathrm{out}}^{i}:=(x_{i}\in\mathrm{Proj}_{x_{i}}(B_{\mathrm{out}}(H))$.
Proof:
(i) Let $\mu_{\mathrm{in}}:=(x\in B_{\mathrm{in}}(H))$. By construction, it holds that $B_{\mathrm{in}}(H)\equiv\mathrm{Proj}_{x_{1}}(B_{\mathrm{in}}(H))\times\cdots\times\mathrm{Proj}_{x_{\nu}}(B_{\mathrm{in}}(H))$, implying that $\mu_{\mathrm{in}}\equiv\wedge_{i=1}^{\nu}\mu_{\mathrm{in}}^{i}$. Since, $B_{\mathrm{in}}(H)\subseteq H$, $\mu_{\mathrm{in}}$ satisfies $\mu$, which leads to $\wedge_{i=1}^{\nu}\mu_{\mathrm{in}}^{i}$ satisfying $\mu$ as needed. (ii) Let $\mu_{\mathrm{out}}:=(x\in B_{\mathrm{out}}(H))$. Similarly to (i), it holds that $\mu_{\mathrm{out}}\equiv\wedge_{i=1}^{\nu}\mu_{\mathrm{out}}^{i}$, since $B_{\mathrm{out}}(H)\equiv\mathrm{Proj}_{x_{1}}(B_{\mathrm{out}}(H))\times\cdots\times\mathrm{Proj}_{x_{\nu}}(B_{\mathrm{out}}(H))$, by construction. Then, we have $\neg\mu_{\mathrm{out}}\equiv\lor_{i=1}^{\nu}\neg\mu_{\mathrm{out}}^{i}$. Since $\wedge_{i=1}^{\nu}\neg\mu_{\mathrm{out}}^{i}$ satisfies $\lor_{i=1}^{\nu}\neg\mu_{\mathrm{out}}^{i}$ and $\neg\mu_{\mathrm{out}}$ satisfies $\neg\mu$, we have that $\wedge_{i=1}^{\nu}\neg\mu_{\mathrm{out}}^{i}$ satisfies $\neg\mu$, completing the proof. ∎
Example 2: Consider formulas $\phi=\square_{[a,b]}\mu$ and $\psi=\square_{[a,b]}\neg\mu$, where $\mu:=(h(x)\geq\bm{0})$ is a predicate, with $h(x)=\begin{bmatrix}c_{1}&c_{2}\end{bmatrix}x+\bm{1}_{4}$, $c_{1}=\begin{bmatrix}-\bm{1}_{2}^{\intercal}&\bm{1}_{2}^{\intercal}\end{bmatrix}^{\intercal}$, $c_{2}=\begin{bmatrix}1&-1&1&-1\end{bmatrix}^{\intercal}$. We wish to convert $\phi$ and $\psi$ into conjunctive forms of two conjuncts (subformulas). First, we define the polytope $H=\{x\mid h(x)\geq\bm{0}\}$, and construct orthotopes $B_{\mathrm{in}}(H)=\{x\mid b_{\mathrm{in}}(x)\geq\bm{0}\}$ and $B_{\mathrm{out}}(H)=\{x\mid b_{\mathrm{out}}(x)\geq\bm{0}\}$, with $b_{\mathrm{in}}(x)=2\begin{bmatrix}\bm{I}_{2}&-\bm{I}_{2}\end{bmatrix}^{\intercal}x+\bm{1}_{4}$ and $b_{\mathrm{out}}(x)=\begin{bmatrix}\bm{I}_{2}&-\bm{I}_{2}\end{bmatrix}^{\intercal}x+\bm{1}_{4}$, respectively, as in Definition 4. Polytope $H$ and orthotopes $B_{\mathrm{in}}(H)$, $B_{\mathrm{out}}(H)$ are depicted in Figure 4. Then, using Lemma 5, we define predicates $\mu_{\mathrm{in}}^{i}:=(x_{i}\in\mathrm{Proj}_{x_{i}}(B_{\mathrm{in}}(H))=(-0.5\leq x_{i}\leq 0.5)$, $\mu_{\mathrm{out}}^{i}:=(x_{i}\in\mathrm{Proj}_{x_{i}}(B_{\mathrm{out}}(H))=(-1\leq x_{i}\leq 1)$, with $i\in\{1,2\}$, allowing us to decompose $\mu$ and $\neg\mu$ as $\mathrm{dec}(\mu,2)=\mu_{\mathrm{in}}^{1}\wedge\mu_{\mathrm{in}}^{2}$ and $\mathrm{dec}(\neg\mu,2)=\neg\mu_{\mathrm{out}}^{1}\wedge\neg\mu_{\mathrm{out}}^{2}$. Finally, $\mathrm{dec}(\phi,2)=\phi^{1}\wedge\phi^{2}$ and $\mathrm{dec}(\psi,2)=\psi^{1}\wedge\psi^{2}$, where $\phi^{i}=\square_{[a,b]}\mu_{\mathrm{in}}^{i}$ and $\psi^{i}=\square_{[a,b]}\neg\mu_{\mathrm{out}}^{i}$, with $i\in\{1,2\}$, are valid decompositions of $\phi$ and $\psi$, respectively.
@@ IV-A Specification Decomposition
Figure 4: Polytopes $H$, $B_{\mathrm{in}}(H)$, $B_{\mathrm{out}}(H)$, for Example 1.
IV-B Heuristic Agent-Specification Filtering
STL is equipped with robustness metrics for assessing the robust satisfaction of a formula [18]. A scalar-valued function $\rho^{\phi}(\bm{x},k)$ of a signal $\bm{x}$ and time $k$ indicates how robustly a signal $\bm{x}$ satisfies a formula $\phi$ at time $k$. The robustness function is defined recursively in [18, Definition 3].
Let $\phi_{k}$ be a newly assigned STL specification at time $k$, and assume that $\mathrm{dec}(\phi_{k},\nu_{k})=\wedge_{j=1}^{\nu_{k}}\phi_{k}^{j}$. In the following, we assume that the total number of agents, denoted by $|\mathcal{I}|$, satisfies $|\mathcal{I}|>\nu_{k}$, otherwise, the following filtering procedure is ignored. Although, $\phi_{k}$ is split into $\nu_{k}$ agent-level sub-specifications, the assignment of $\phi_{k}^{j}$ to a specific agent $i\in\mathcal{I}$ is not determined by $\mathrm{dec}(\phi_{k},\nu_{k})$. In fact, each sub-specification, $\phi_{k}^{j}$, may be assigned to any agent $i\in\mathcal{I}$ resulting in $\nu_{k}|\mathcal{I}|$ agent-specification pairs. To reduce the number of possible pairs, we utilize the following heuristic approach: Based on the quantitative semantics of STL introduced earlier, we compute the robustness function of $\phi_{k}^{j}$ over the $i$th agent’s trajectory. This metric indicates roughly the distance between the trajectory of agent $i$ and the space of signals that satisfy $\phi_{k}^{j}$ [25], thereby facilitating a preliminary agent-specification assessment. We approximate the $i$th agent’s trajectory by $\bm{z}^{i,k-1}=\{z_{0}^{i,k-1},z_{1}^{i,k-1},\ldots\}$, $i\in\mathcal{I}$, that is, the signal associated with the (deterministic) nominal dynamics of agent $i$, (4a), computed at time $k-1$, which is available at time $k$. Specifically, for each $j\in\{1,\ldots,\nu_{k}\}$, and $i\in\mathcal{I}$, we efficiently calculate $\rho^{\phi_{k}^{j}}(\bm{z}^{i,k-1},k)$, which requires negligible computational cost. Subsequently, for each $\phi_{k}^{j}$, we collect the largest $\nu_{k}$ values of $\rho^{\phi_{k}^{j}}(\bm{z}^{i,k-1},k)$ among the $|\mathcal{I}|$ options, resulting in $(\nu_{k})^{2}$ agent-specification pairs. Should $v_{k}<<|\mathcal{I}|$, the result is a significant reduction.
IV-C Task Allocation Algorithm
To determine an agent-specification assignment, we solve the following binary optimization problem, given by
$\min_{\bm{\lambda}}\ \sum_{(i,j)\in\mathbb{L}}$ $\lambda_{i,j}r^{i}_{\phi_{j},mpc},$ (12a)
[…]
$\sum_{j\in\{1,\cdots,\nu\}}\lambda_{i,j}$ $\leq 1,\ \forall i\in\mathcal{I},$ (12c)
@@ IV-C Task Allocation Algorithm
IV-D Implementation & Theoretical Analysis
@@
1:Given: System (1) initialized at $x^{i}(0)$ for all $i\in\mathcal{I}$
2:Get normalized dynamics of (1) for all $i\in\mathcal{I}$,
3:for $k\in\{0,\cdots,N-1\}$ do
4:     if no new spec. then set $\mathbb{L}_{opt}=\emptyset$ go to line 16
5:     Task Allocation:
6:     ($\psi$, $r_{\psi,\max}$) $\leftarrow$ load new specification
7:     Decompose into ($\psi^{j}$, $r_{\psi^{j},\max}$), $j\in\{1,\cdots,\nu\}$
8:     Filter pairs to obtain $\mathbb{L}\subset\mathcal{I}\times\{1,\cdots,\nu\}$
9:     for $(i,j)\in\mathbb{L}$ do
10:         Solve (9), get $r^{i}_{\psi^{j},mpc}$ and control strategy $\bm{f}^{(i,j)}$
11:         if no solution then $\mathbb{L}=\mathbb{L}\setminus\{(i,j)\}$
12:     end for
13:     Solve (12) to get definitive assignment list $\mathbb{L}_{F}$
14:     Control Strategy Update:
15:     $\forall(i,j)\in\mathbb{L}_{opt}$, set $\bm{f}^{i,k}=\bm{f}^{(i,j)}$leftarrow update
16:     for $i\in\mathcal{I}\setminus\mathrm{Proj}_{\mathcal{I}}(\mathbb{L}_{F})$ do
17:         Solve (9) to get control strategy $\bm{f}^{i}_{meas}$
18:         if solution then set $\bm{f}^{i,k}=\bm{f}^{i}_{meas}$leftarrow update
19:         else set $\bm{f}^{i,k}=\bm{f}^{i,k-1}$leftarrow no update
20:     end for
21:     Implement $\bm{f}^{\mathcal{I},k}$ and measure $x^{i}(k+1),\ \forall i\in\mathcal{I}$
22:end for
@@
It is sufficient to prove that the existence of $\bm{f}^{i,k}$ implies the existence of $\bm{f}^{i,k+1}$ for each agent $i\in\mathcal{I}$. Notice that should an agent either accept a new sub-specification or measurement only, an updated control strategy is obtained, respectively, as $\bm{f}^{i,k+1}=\bm{f}^{(i,j)}$ or $\bm{f}^{i,k+1}=\bm{f}^{i}_{meas}$. Should neither be available, the previous control strategy is again implemented, $\bm{f}^{i,k+1}=\bm{f}^{i,k}$, completing the proof. ∎
For each specifications $\phi$ accepted at time $k$, by subgroup $I\subseteq\mathcal{I}$, we need that $\mathbb{P}_{\bm{f}}(\phi,k)\geq 1-r_{\phi,\max}$. Here, it is sufficient to show that $\forall i\in I$, $\mathbb{P}_{\bm{f}}(\phi^{i},k)\geq 1-r_{\phi^{i},\max}$. The latter, in relation to optimization problem (9), is discussed within open-loop and closed-loop implementation in our previous work [6]. In said paper, we established the validity of open-loop implementation and suggested that closed-loop implementation follows naturally from the convex unimodality of the disturbance.
@@
Note that the local risk value $r_{\phi^{j},mpc}$ compared to the average maximal risk $r_{\phi^{j},\max}$ is conservative, $r_{\phi^{j},mpc}\geq r_{\phi^{j},\max}$, due to a union-bound argument used in proving Prop. 3. Also, the obtained task allocation may not always be optimal or feasible due to our heuristic choice of filtering, even if an allocation exists.
V Case Study

### v2

@@ III-D Control Synthesis and Local Risk Value
In this subsection, we compute the local risk value and synthesize a controller for a given agent-specification pair utilizing previous work on risk-aware MPC for single-agent systems [15]. We refer the reader to [15] for more details.
We assume a new specification $\phi$ is provided at time $k$ together with maximal risk $r_{\phi,\max}$. For any specifications $\psi$, we denote with $k_{\psi}$ the time at which the specification was assigned. We denote with $\mathcal{P}^{k}$ the set of all previously accepted specifications $\psi$ with $k_{\psi}<k$, together with the newly assigned specification $\phi$, i.e., $\phi\in\mathcal{P}^{k}$. The objective is to synthesize a controller $\bm{f}$ such that $\varphi:=\wedge_{\psi\in\mathcal{P}^{k}}\lozenge_{[k_{\psi},k_{\psi}]}\psi$ is satisfied in accordance to the maximal risk of each conjunction element, i.e., $\mathbb{P}_{\bm{f}}(\psi,k_{\psi})\geq 1-r_{\psi,\max},\ \forall\psi\in\mathcal{P}^{k}$. To accomplish the objective, we make the following assumption.
Assumption 2
Distribution $\mathcal{Q}_{w}$ has zero mean and $A_{K}\Sigma_{\infty}A_{K}^{T}+\text{var}(\mathcal{Q}_{w})=\Sigma_{\infty}$ has solution $\Sigma_{\infty}=I_{n}$
Note that the above assumption is not restrictive as a simple coordinate transformation $y=\Sigma_{\infty}^{-\frac{1}{2}}x$ will ensure the second part of the assumption is satisfied. We call the transformed dynamics the normalized dynamics, and in the remainder, we maintain our notations as if the dynamics are normalized.
We are interested in implementing a tube-based MPC algorithm that at each time instance $k\in[0,N]$ recomputes the optimal nominal trajectory $\bm{z}_{[k,N]}$, the optimal nominal input $\bm{v}_{[k,N-1]}$, and the optimal risk levels $\bm{r}_{[k,N]}$. For this, we assume given $x(k)$ together with $\bm{z}_{[0,k]}$, $\bm{v}_{[0,k-1]}$, $\bm{\rho}_{[0,k]}$ and $\bm{r}_{[0,k]}$ computed at time $k-1$. Furthermore, for all $\psi\in\mathcal{P}^{k}$, we let $\mathcal{H}_{\psi}^{k_{\psi}}\subseteq[0,N]$ without loss of generality. Here, the active horizon $\mathcal{H}_{\phi}^{k}$ is the set of all time instances needed to evaluate $\bm{x}_{k}\vDash\phi$ (see [15, Definition 4]). It should be noted that any active horizon consists of only finite elements.
@@
where $\Xi:=\{\bm{z}_{[k,N]},\bm{v}_{[k,N-1]},\bm{r}_{[k+1,N]},\bm{\rho}_{[k+1,N]},s(k)\}$, $J$ is either a linear or quadratic cost function, the constraints (5d)-(5e) are obtained from $\varphi=\wedge_{\psi\in\mathcal{P}^{k}}\lozenge_{[k_{\psi},k_{\psi}]}\psi$ and the constraints (5f)-(5g) are obtained from [15, Thm. 4]. From the above tube-based MPC, an (updated) control strategy is obtained by $\bm{v}_{[k,N-1]}$, and a local risk value $r_{\phi,mpc}$ is obtained by the right-hand side of (5g), i.e.,
@@
A solution to (5), by [15, Thm. 4], ensures that for all $\psi\in\mathcal{P}^{k}$, $\mathbb{P}_{\bm{f}}(\psi,k_{\psi})\geq 1-r_{\psi,\max}$. Should there exist no solution to problem (5), the newly provided agent-specification pair is deemed infeasible. In such a case, we will update the control strategy based only on measurement $x(k)$, as per constraint (5b). This can be achieved by considering the MPC problem (5) with $\mathcal{P}^{k}=\mathcal{P}^{k}\setminus\{\phi\}$. However, should the measurement also fail to produce a controller, the previously computed control strategy will be implemented instead.
III-E Task Allocation Algorithm
To determine an agent-specification assignment, we solve the following binary optimization problem given by
$\textstyle\min_{\bm{\lambda}}\ \sum_{(i,j)\in\mathbb{L}}$ $\lambda_{i,j}r^{i}_{\phi_{j},mpc},$ (7a)
@@ III-E Task Allocation Algorithm
$\textstyle\sum_{j\in\{1,\cdots,\nu\}}\lambda_{i,j}$ $\leq 1,\ \forall i\in\mathcal{I},$ (7c)
@@
III-F Implementation & Theoretical Analysis
@@
1:Given: Normalized system (1) with $x^{i}(0),\ \forall i\in\mathcal{I}$
2:for $k\in\{0,\cdots,N-1\}$ do
3:     if no new spec. then set $\mathbb{L}_{opt}=\emptyset$ go to line 15
4:     Task Allocation:
5:     ($\psi$, $r_{\psi,\max}$) $\leftarrow$ load new specification
6:     Decompose into ($\psi^{j}$, $r_{\psi^{j},\max}$), $j\in\{1,\cdots,\nu\}$
7:     Filter pairs to obtain $\mathbb{L}\subset\mathcal{I}\times\{1,\cdots,\nu\}$
8:     for $(i,j)\in\mathbb{L}$ do
9:         Solve (5), get $r^{i}_{\psi^{j},mpc}$ and control strategy $\bm{f}^{(i,j)}$
10:         if no solution then $\mathbb{L}=\mathbb{L}\setminus\{(i,j)\}$
11:     end for
12:     Solve (7) to get definitive assignment list $\mathbb{L}_{opt}$
13:     Control Strategy Update:
14:     $\forall(i,j)\in\mathbb{L}_{opt}$, set $\bm{f}^{i,k}=\bm{f}^{(i,j)}$leftarrow update
15:     for $i\in\mathcal{I}\setminus\mathrm{Proj}_{\mathcal{I}}(\mathbb{L}_{opt})$ do
16:         Solve (5) to get control strategy $\bm{f}^{i}_{meas}$
17:         if solution then set $\bm{f}^{i,k}=\bm{f}^{i}_{meas}$leftarrow update
18:         else set $\bm{f}^{i,k}=\bm{f}^{i,k-1}$leftarrow no update
19:     end for
20:     Implement $\bm{f}^{\mathcal{I},k}$ and measure $x^{i}(k+1),\ \forall i\in\mathcal{I}$
21:end for
@@
It is sufficient to prove that the existence of $\bm{f}^{i,k}$ implies the existence of $\bm{f}^{i,k+1}$ for each agent $i\in\mathcal{I}$. Notice that if an agent either accepts a new sub-specification or measurement only, an updated control strategy is obtained, respectively, as $\bm{f}^{i,k+1}=\bm{f}^{(i,j)}$ or $\bm{f}^{i,k+1}=\bm{f}^{i}_{meas}$. If neither is available, the previous control strategy is again implemented, $\bm{f}^{i,k+1}=\bm{f}^{i,k}$, completing the proof. ∎
For each specifications $\phi$ accepted at time $k$, by subgroup $I\subseteq\mathcal{I}$, we need that $\mathbb{P}_{\bm{f}}(\phi,k)\geq 1-r_{\phi,\max}$. Here, it is sufficient to show that $\forall i\in I$, $\mathbb{P}_{\bm{f}}(\phi^{i},k)\geq 1-r_{\phi^{i},\max}$. This, in relation to optimization problem (5), is discussed within open-loop and closed-loop implementation in our previous work [15]. In said paper, we established the validity of open-loop implementation and suggested that closed-loop implementation follows naturally from the convex unimodality of the disturbance.
@@
Note that the local risk value $r_{\phi^{j},mpc}$ compared to the average maximal risk $r_{\phi^{j},\max}$ is conservative, $r_{\phi^{j},mpc}\geq r_{\phi^{j},\max}$, due to a union-bound argument used in proving [15, Thm. 4]. Also, the obtained task allocation may not always be optimal or feasible due to our heuristic choice of filtering, even if an allocation exists.
IV Case Study
We validate our proposed method using a multi-shuttle routing case. In a tourist attraction spot (Fig. 3), four shuttle buses of differing speeds (A-D) should pick up tourists from gathering points (GP) and deliver them to an unloading point (ULP) while avoiding two buildings (B). The shuttles initiate from their respective terminals (T) and should return when required. Each task should be assigned to a most probable shuttle to fetch the tourists. Thus, new routing may be assigned anytime, rendering a dynamic task allocation problem. The technical details can be found in the Appendix.
Figure 3: The shuttle trajectories.
After decomposing, filtering, control synthesis, and auctioning, we obtain the results in Fig. 3. The logging information of the experiment indicates that: 1) at time $k=2$, buses ‘A’, ‘B’, and ‘D’ are assigned with GP-II, GP-IV, and GP-I, respectively; 2) at time $k=8$, buses ‘A’ and ‘D’ are assigned with GP-II and GP-III, respectively; 3) at time $k=15$, buses ‘B’, ‘C’, and ‘D’ have accepted the return specification, while shuttle ‘A’ has rejected it. Bus ‘C’ is not assigned any task throughout the experiment and, as a result, stays at its terminal. Since buses ‘A’ and ‘D’ accept additional tasks at $k=8$, our method allows agents to adjust their behaviours according to newly assigned tasks. As mentioned, shuttle ‘A’ has rejected the return task at time $k=15$. This is due to the incapability to move long distances considering its small speed limit. Additionally, all buses are confined within the ‘BOX’ region and kept outside the buildings, implying the compatibility of our method to explicitly assigned specifications. Part of the trajectory of shuttle ‘D’ intersects with B-2 due to naive linear interpolation. In practice, nonlinear interpolation can guarantee critical safety, which may generate curved trajectories when navigating around the building.
Remark 2
V Conclusion
This paper considers real-time task allocation and control synthesis for stochastic heterogeneous linear multi-agent systems with probabilistic satisfaction of STL specifications. The proposed method relies on the dynamic allocation of decomposed specifications to individual agents using risk measures via auctioning. The risk measures are computed using tube-based MPC on the agent level during the control synthesis. These two approaches provide a novel perspective for coordination control of multi-agent systems in dynamic environments. In future work, we will improve the risk quantification for less conservative control solutions.
Appendix
V-A Problem Statement: Illustrative example
We consider an example (see Fig. 4) to further illustrate the problem and approach covered within this paper.
V-B Specification Decomposition: Illustrative example
We consider an example (see Fig. 5) to illustrate the spe-cification decomposition approach covered within this paper.
Example 2: Consider formulas $\phi=\square_{[a,b]}\mu$ and $\psi=\square_{[a,b]}\neg\mu$, where $\mu:=(h(x)\geq\bm{0})$ is a predicate, with $h(x)=\begin{bmatrix}c_{1}&c_{2}\end{bmatrix}x+\bm{1}_{4}$, $c_{1}=\begin{bmatrix}-\bm{1}_{2}^{\intercal}&\bm{1}_{2}^{\intercal}\end{bmatrix}^{\intercal}$, $c_{2}=\begin{bmatrix}1&-1&1&-1\end{bmatrix}^{\intercal}$. We wish to convert $\phi$ and $\psi$ into conjunctive forms of two conjuncts (subformulas). First, we define the polytope $H=\{x\mid h(x)\geq\bm{0}\}$, and construct orthotopes $B_{\mathrm{in}}(H)=\{x\mid b_{\mathrm{in}}(x)\geq\bm{0}\}$ and $B_{\mathrm{out}}(H)=\{x\mid b_{\mathrm{out}}(x)\geq\bm{0}\}$, with $b_{\mathrm{in}}(x)=2\begin{bmatrix}\bm{I}_{2}&-\bm{I}_{2}\end{bmatrix}^{\intercal}x+\bm{1}_{4}$ and $b_{\mathrm{out}}(x)=\begin{bmatrix}\bm{I}_{2}&-\bm{I}_{2}\end{bmatrix}^{\intercal}x+\bm{1}_{4}$, respectively, as in Definition 2. Polytope $H$ and orthotopes $B_{\mathrm{in}}(H)$, $B_{\mathrm{out}}(H)$ are depicted in Fig. 5. Then, using Lemma 3, we define predicates $\mu_{\mathrm{in}}^{i}:=(x_{i}\in\mathrm{Proj}_{\mathbb{X}^{i}}(B_{\mathrm{in}}(H))=(-0.5\leq x_{i}\leq 0.5)$, $\mu_{\mathrm{out}}^{i}:=(x_{i}\in\mathrm{Proj}_{\mathbb{X}^{i}}(B_{\mathrm{out}}(H))=(-1\leq x_{i}\leq 1)$, with $i\in\{1,2\}$, allowing us to decompose $\mu$ and $\neg\mu$ as $\mathrm{dec}(\mu,2)=\mu_{\mathrm{in}}^{1}\wedge\mu_{\mathrm{in}}^{2}$ and $\mathrm{dec}(\neg\mu,2)=\neg\mu_{\mathrm{out}}^{1}\wedge\neg\mu_{\mathrm{out}}^{2}$. Finally, $\mathrm{dec}(\phi,2)=\phi^{1}\wedge\phi^{2}$ and $\mathrm{dec}(\psi,2)=\psi^{1}\wedge\psi^{2}$, where $\phi^{i}=\square_{[a,b]}\mu_{\mathrm{in}}^{i}$ and $\psi^{i}=\square_{[a,b]}\neg\mu_{\mathrm{out}}^{i}$, with $i\in\{1,2\}$, are valid decompositions of $\phi$ and $\psi$, respectively.